OpenStudy (anonymous):

When 40.0 mL of 1.00 M H2SO4 is added to 80.0 mL of 1.00 M NaOH at 20.00°C in a coffee cup calorimeter, the temperature of the aqueous solution increases to 29.20°C. If the mass of the solution is 120.0 g and the specific heat of the calorimeter and solution is 4.184 J/g • °C, how much heat is given off in the reaction

OpenStudy (abmon98):

q=m*c*Delta T Q=120*4.184*(29.20-20)